40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 21 12:25:46 2019, gautam, Update, ALS, DFD electronics checkout ALSnoiseAnomaly.pdfDFDcalib.pdfspectra.pdfPTgains.pdf
    Reply  Tue Oct 22 15:32:15 2019, gautam, Update, ALS, EX uPDH electronics checkout uPDH_X_OLTFs.pdf
       Reply  Wed Oct 23 10:23:26 2019, rana, Update, ALS, EX uPDH electronics checkout 
          Reply  Wed Oct 23 11:11:01 2019, gautam, Update, ALS, EX uPDH electronics checkout 
    Reply  Fri Oct 25 01:04:49 2019, gautam, Update, ALS, ALS electronics chain was saturating 
       Reply  Mon Oct 28 23:20:11 2019, gautam, Update, ALS, ALS power budget 
Message ID: 14993     Entry time: Fri Oct 25 01:04:49 2019     In reply to: 14981     Reply to this: 14995
Author: gautam 
Type: Update 
Category: ALS 
Subject: ALS electronics chain was saturating 

[Koji, gautam]

Summary:

We think we got to the bottom of this issue today. The RF signal level going into the demod board is too high. This electronics chain needs some careful gain reallocation.

Details:

I was demonstrating to Koji a strange feature I had noticed in the ALS control, whereby when applying a CARM offset to detune the arms, the two arms seemed to respond differently (based on the transmission levels). This kind of CARM-->DARM coupling seemed strange to me. Anyway, I also noticed that the EPICS indicators on the ALS MEDM screen suggested ADC saturations were going on. I had never really looked at the fast time series of the inputs to the phase tracker servos, but these showed saturating behavior on ndscope traces. I went to the LSC rack and measured these on a scope, indeed, they were ~20V pp.

The output of the BeatMouth PDs are going to a ZHL-3A amplifier - we should consider replacing these with lower gain amplifiers, e.g. the Teledyne AP1053. This is relegated to a daytime task.

Other findings tonight:

While working on the PSL table, I somehow put the IMC FSS into a bad state, reminiscent of this behavior. Seems like this is linked to some flaky connection on the PSL table. One candidate is the unstable attachment of the Pomona box between the NPRO PZT and the FSS output - we should install a short BNC cable between these to avoid the lever arm situation we have right now.

ELOG V3.1.3-